# Look Before You Leap: Problem Elaboration Prompting Improves Mathematical Reasoning in Large Language Models

### v1

@@ 4.1.   Setup
We evaluate PEP on four elementary datasets, with a focus on mathematical reasoning: (1) SingleEq (Koncel-Kedziorski et al., 2015), (2) GSM8k (Cobbe et al., 2021), (3) SVAMP (Patel et al., 2021), (4) AQuA (Ling et al., 2017); and five challenging subsets from MMLU-STEM Hendrycks et al. (2020), denoted by (5) MMLU-MATH. Furthermore, we investigate the distraction problem using GSMIC (Shi et al., 2023): we randomly sampled 500 examples separately for 2-step problems and m-step problems, denoted by (6) GSMIC-1k. Statistical information are presented in Table 3.
@@
Model Prompt PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT (Kojima et al., 2022) 87.40 50.19 38.19 72.1 61.97
90.35 60.12 40.55 75.7 66.68
(+2.95) (+9.93) (+2.36) (+3.6) (+4.71)
PoT (Chen et al., 2022) 95.67 63.84 29.13 78.3 66.74
96.46 64.75 29.53 81.3 68.01
GPT-3.5 text-davinci-003 (+0.79) (+0.91) (+0.40) (+3.0) (+1.27)
CoT (Kojima et al., 2022) 91.93 80.44 55.12 78.8 76.57
91.93 78.92 57.09 79.8 76.94
(+0.0) (-1.52) (+1.97) (+1.0) (+0.37)
PoT (Chen et al., 2022) 97.05 75.82 36.61 81.6 72.77
96.06 76.57 37.80 84.1 73.63
ChatGPT gpt-3.5- turbo-0301 (-0.99) (+0.75) (+1.19) (+2.5) (+0.86)
Table 4: Accuracies (x100) of our approach (PEP) against zero-shot CoT and zero-shot PoT. All results are with greedy decoding (GD), i.e., the temperature is set to zero. PEP demonstrate an overall improvements compared to baselines, especially with davinci on GSM8k ($50.19\%\to 60.12\%$).
@@
Model Prompt PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT+SC (Wang et al., 2022c) 87.99 52.31 39.76 71.7 62.94
@@
GPT-3.5 text-davinci-003 (+1.58) (+8.80) (+0.0) (1.90) (+3.07)
CoT+SC (Wang et al., 2022c) 93.50 88.70 69.29 84.6 84.02
93.50 90.98 68.11 86.2 84.70
ChatGPT gpt-3.5- turbo-0301 (+0.0) (+2.28) (-1.18) (+1.6) (+0.68)
Table 5: Solve rates (x100) of our approach (PEP) against standard prompting (zero-shot CoT). All results are with self-consistency (SC), where temperature T=0.7 and reasoning paths are sampled for 20 times.
Baselines & Prompts.
We compare PEP against (1) zero-shot CoT (Kojima et al., 2022) for textual reasoning, and (2) zero-shot PoT (Chen et al., 2022) for code-based reasoning. We also compare (3) Irr-Inst. that Shi et al. (2023) suggests to mitigate distraction problem. We added some conjunctions and adjusted the template of PoT to ensure the format is valid. All prompts are presented in the Appendix in detail.
Language Models & Decoding
We conduct our main experiments on (1) the “text-davinci-003” model  (denoted as davinci hereafter) and (2) the “gpt-3.5-turbo-0301” model (denoted as turbo hereafter) (Brown et al., 2020; Ouyang et al., 2022b; OpenAI, 2023). Both models are employed via OpenAI API. We evaluate the performance with both greedy decoding and self-consistency decoding (SC) (Wang et al., 2022c) to ensure the reproducibility of experiments. There may be variations in the results obtained from other versions or updated models.
4.2.   Main Results
The main results are presented in Table 4 and Table 5. Considering the token cost, we only compared PEP with CoT using self-consistency. Overall, our proposed PEP performs better compared to the vanilla CoT and PoT, achieving SOTAs with turbo (90.98% on GSM8k and 86.2% on SVAMP). We also evaluate PEP against vanilla CoT on challenging MMLU-MATH, which includes five more complex mathematical tasks, as shown in Fig. 3. PEP presents consistently improvements for most subtopics with both davinci ($\xrightarrow{}$44.2%) and turbo ($\xrightarrow{}$56.0%).
Figure 3: Performances on MMLU-Math. All results are generated with greedy search. Although there is a decline for the HighSchoolMath subtopic, PEP exhibits an overall enhancement compared to the vanilla CoT across the complex subtopics.
Micro Accuracy Macro Accuracy
Method (turbo-0301) 2 steps >2 steps Overall Norm 2 steps >2 steps Overall Norm
CoT 83.2 74.2 78.7 85.42 83.26 72.05 77.65 84.41
PEP 83.6 75.6 79.6 86.4 83.11 73.99 78.55 85.38
(+0.4) (+1.4) (+0.9) (+0.98) (-0.15) (+1.95) (+0.90) (+0.98)
@@ 4.2.   Main Results
CoT 85.6 80.8 83.2 90.3 86.28 79.19 82.74 89.93
PEP 87.4 84.0 85.7 93.02 86.63 83.83 85.23 92.65
(+1.8) (+3.2) (+2.5) (+2.71) (+0.35) (+4.65) (+2.5) (+2.72)
Table 6: Micro and macro accuracies (×100) on GSMIC-1k. We follow the metrics and instruction suggested in (Shi et al., 2023). Norm is the overall accuracy normalized by the scores of base problems, measuring how a method is affected by the distractors. All results are generated with greedy decoding.
Specifically, when working with davinci, which may be less capable in problem recognition compared to turbo, PEP shows significant improvement and outperforms the baselines by a substantial margin (9.93% and 8.80% on GSM8k for two decoding settings separately). Under greedy decoding, PEP exhibits an overall superiority to both CoT and PoT. The enhancement for davinci is more pronounced with CoT, whereas for turbo, it is PoT. This could suggest that PEP does not have a particular reliance on the manner of downstream reasoning. In addition, PEP with turbo shows a decrease on GSM8k under GD, but performs better with SC. In Table 5, PEP achieves the same accuracy as CoT for some cases and performs worse on AQuA. To further analyze the performance of PEP with SC, we provide a break-down analysis in the Sec. 4.4.
4.3.   Distraction Problem
LLMs may encounter difficulties when handling ill-formed problems. One particular challenge is known as the distraction problem (Shi et al., 2023), which means when adding irrelevant sentences to a given problem, the model can be easily distracted and further generate erroneous results. These additional sentences can be completely irrelevant or relevant to the problem but have no impact on the reasoning. An example is in Table 7, comparing the performances of CoT and PEP when dealing with the distracting problem.
In this study, we investigated PEP on distraction problem based on GSMIC dataset (Shi et al., 2023), which is constructed from a base subset of GSM8k (Cobbe et al., 2021). The irrelevant sentences are randomly added according to certain templates and pre-designed distracting factors. Two metrics are used in our experiments: (1) micro-accuracy: averaged accuracy per example, and (2) macro accuracy: averaged accuracy per base problem. The main results are shown in Table 6.
Figure 4: Break-down analysis of the number of self-consistent sampling paths. Comparing with the results in Table 5, we can find PEP presents consistently better performances than the standard CoT for most cases.
Q: The average number of fruits per basket in five baskets is 25. Ada bought 20 tomatoes from the grocery store. If basket A has 15 apples, B has 30 mangoes, C has 20 peaches, D has 25 pears and E has some bananas, how many bananas are in basket E?
Standard CoT (turbo with greedy decoding):
A: 1. Total number of fruits: 25 x 5 = 125. 2. Add the tomatoes to the total number of fruits: 125 + 20 = 145. 3. Subtract the number of fruits in baskets A, B, C, and D from the total number of fruits: Bananas in basket E = 145 - 90 = 55.  (Wrong: distracted.)
PEP + CoT (turbo with greedy decoding):
(+Elaboration; Interrogative):
Segment 1: What is the average number of fruits per basket in five baskets?
Segment 2: How many tomatoes did Ada buy from the grocery store?
Segment 3: What is the fruit distribution in each basket?
Segment 4: How many bananas are in basket E?
A: Segment 1: The average number is 25.
Segment 2: Ada bought 20 tomatoes.
Segment 3: A: 15 apples, B: 30 mangoes, C: 20 peaches, D: 25 pears, E: some bananas.
Segment 4: To find out how many bananas are in basket E, we need to use the information given in Segment 1 and Segment 3.
Total number is 25 x 5 = 125. In A, B, C, and D is 15 + 30 + 20 + 25 = 90. Therefore, the number of fruits in basket E is 125 - 90 = 35.  (Correct: concentrated.)
Table 7: When adding irrelevant sentence (cyan) to the problem, CoT is easily distracted (red), while PEP is more focused (blue). We omitted and rephrased certain sentences for clarity.
PEP can help alleviate the distraction problem.
In Table 6, PEP outperforms CoT for both metrics, demonstrating better robustness to distraction problem. We can further observe that PEP performs better on m-step problems than 2-step tasks, indicating a more significant improvement where the premises are intricately coupled with irrelevant factors. The macro accuracy of PEP declines for 2-step problems, indicating the performance of PEP may vary for different base problems.
PEP performs better when prompted with prior knowledge.
When the model is consciously prompted to ignore irrelevant content for the given problem, referred to as Irr-Inst. (Shi et al., 2023), we observed significant improvements in both CoT and PEP. However, PEP exhibits more substantial and consistent enhancements compared to the CoT, with a larger margin ($0.98\%\to 2.71\%$). It is promising that PEP may also be effective for other types of ill-formed problems with prior knowledge.
4.4.   Ablation Study
Method AQuA GSM8k GSMIC-1k
CoT 55.12 80.44 78.7
PEP 57.09 78.92 79.6
- Dec 53.15 78.17 76.7
- Elu 62.6 78.70 78.6
EtD 54.33 76.27 77.1
Table 8: Ablation Study. Dec and Elu indicate prompting the model to only perform decomposition and perform elucidation, respectively. EtD means prompting the model to elucidate then decompose.
In this section, we conducted ablation studies to verify the criticality of PEP principles, performances of different SC sampling settings, and the effect of diverse distracting factors.
Micro Accuracy Macro Accuracy
Topic Role Overlap Num. Range Topic Role Overlap Num. Range
Method (turbo) In Off Yes No In Out In Off Yes No In Out
CoT 68.58 87.71 74.95 81.94 78.43 78.97 66.86 87.15 76.49 81.72 78.68 78.44
PEP 73.67 84.88 79.27 79.89 78.43 80.75 69.9 86.11 78.23 80.45 78.01 80.04
(+5.1) (-2.84) (+4.32) (-2.05) (+0.0) (+1.79) (+3.04) (-1.04) (+1.73) (-1.27) (-0.67) (+1.6)
+ Irr-Inst. (Shi et al., 2023)
CoT 75.58 89.98 83.8 82.68 83.27 83.13 74.75 89.84 84.46 82.24 82.48 83.4
PEP 79.19 91.49 85.53 85.85 83.06 88.29 77.47 91.76 84.61 86.16 82.61 86.98
(+3.61) (+1.51) (+1.73) (+3.17) (-0.2) (+5.16) (+2.72) (+1.92) (+0.14) (+3.92) (+0.13) (+3.58)
@@ 4.4.   Ablation Study
Table 9: The breakdown accuracies (×100) w.r.t. the added irrelevant sentence factors. Lower accuracy suggests the model is more sensitive to irrelevant contexts of that type.
Example 1 (PEP misleads implicit meaning)
Example 2 (PEP discards important information)
Q: An interior design firm offers installation for \$129.00. It includes hanging 4 mirrors, 2 shelves, 1 chandelier, and 10 pictures. They will install additional items for an extra \$15.00 per item. Angela has 6 mirrors and 2 chandeliers and 20 pictures that she needs installed/hung. How much will this cost her?
Q: Mrs. Harrington bought 12 boxes of oranges at the market. She gave her mom and her sister 2 boxes of oranges each. Then she kept 1/4 of the oranges and sold the rest. How many oranges did Mrs. Harrington sell if each box contains 20 oranges?
PEP + CoT; Mixed:
PEP + CoT; Q/A pairs:
S1: What is the base installation fee for the interior design firm? \$129.00. S2: What is included in the base installation fee? 4 mirrors, 2 shelves, 1 chandelier, and 10 pictures. S3: How much does the interior design firm charge for each additional item installed? Extra \$15.00 per item installed. S4: How many additional items does Angela need installed? Angela needs 6 mirrors, 2 chandeliers, and 20 pictures installed, which is a total of 28 additional items.
Solution: The cost of the additional items is 28 items x \$15.00 per item = \$420.00. Therefore, the total cost for Angela’s installation is \$129.00 + \$420.00 = \$549.00.
S1: How many boxes of oranges did Mrs. Harrington buy at the market? 12 boxes. S2: How many boxes of oranges did Mrs. Harrington give to her mom and her sister? 2 boxes. S3: How many boxes of oranges did Mrs. Harrington keep for herself? Mrs. Harrington kept 1/4 of the oranges.S4: How many boxes of oranges did Mrs. Harrington sell? The rest.
Solution: 1. Calculate the total number of oranges Mrs. Harrington bought: 12x20=240. 2. Calculate the total number of oranges she gave to her mom and her sister: 2x2=4, 4x20=80. 3. Total number of oranges kept for herself: 1/4x240=60. 4. Calculate the total number of oranges sold: 240-80-60=100.
Table 10: Examples for Error Analysis using turbo (GD). When pre-processing the problems, PEP may changes the underline meaning and discarding information (red) of the original sentence (cyan).
@@
We verify two components proposed in PEP, namely (1) Dec: decomposing only and (2) Elu: elucidating only. The results are presented in Table 8. It can be observed that both Dec and Elu are crucial for PEP. Besides, Elu outperforms both CoT and PEP on AQuA, indicating that the effectiveness of PEP’s components may vary depending on the context. For instance, the problem may be challenging but have already been composed of simple segments. Additionally, we evaluated the prompt of elucidating the problem before decomposition (denoted as EtD). EtD consistently performs worse than PEP and CoT. This may be because elucidation first increases the difficulty of decomposition.
Break-down analysis of different sampled reason paths.
To further investigate PEP and CoT with self-consistent voting, we conducted a break-down analysis with different numbers of reasoning paths (as shown in Fig. 4). Our results show that, for both davinci and turbo, PEP outperforms CoT in most cases. However, increasing the path number does not always lead to improved performances.
Breakdown analysis of distracting factors.
We analyzed the performance of CoT and PEP using different distracting factors, shown in Table 9. It can be observed that our proposed PEP performs better in handling of (1) in-topic sentence with, (2) role name overlap, and (3) out-range numbers, indicating the effectiveness of PEP in recognizing and parsing the problem context. PEP may be confused by the completely irrelevant context. When prompted with prior knowledge (Irr-Inst.) to ignore irrelevant sentences in the input context, PEP consistently outperforms standard CoT by a larger margin on most distracting factors.
4.5.   Error Analysis

### v2

@@ 4.1 Setup
We evaluate PEP on four elementary datasets, with a focus on mathematical reasoning: (1) SingleEq (Koncel-Kedziorski et al., 2015), (2) GSM8k (Cobbe et al., 2021), (3) SVAMP (Patel et al., 2021), (4) AQuA (Ling et al., 2017). Furthermore, we investigate the distraction problem using GSMIC (Shi et al., 2023): we randomly sampled 500 examples separately for 2-step problems and m-step problems, denoted by (5) GSMIC-1k. The details are shown in Table 2.
Baselines & Prompts.
We evaluate PEP with two elementary answer types: (1) Chain-of-Thoughts (CoT(Kojima et al., 2022) for textual reasoning, and (2) Program-of-Thoughts (PoT(Chen et al., 2022) for code-based reasoning. Three problem-related methods are compared to PEP: (3) Least-to-Most (L2M(Zhou et al., 2022a) , (4) Plan-and-Solve (PaS)  (Wang et al., 2023b) and (5) Self-ask (Press et al., 2022). To investigate the distraction problem, we also adopt (6) Irr-Inst. suggested by Shi et al. (2023). All instructions and exemplars can be found in the appendix B.
Language Models & Decoding
We conduct our main experiments on four open-source LMs (1) two LLama2 models (Touvron et al., 2023b) and (2) two Mistral models (Jiang et al., 2023a; 2024). We also employ two GPT models (3) “text-davinci-003” (davinci) and (4) recent released “gpt-3.5-turbo-0125” (turbo(Brown et al., 2020; Ouyang et al., 2022b; OpenAI, 2023). We evaluate the performance with both greedy decoding and self-consistency decoding (SC) (Wang et al., 2022c) to ensure the reproducibility of experiments.
4.2 Main Results
@@
Model AnsType PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT 73.82 29.72 21.26 48.1 (+0.33)
LLama2 7B-hf-chat 71.87 27.32 23.62 51.4
CoT 80.91 41.47 24.41 59.7 (+0.68)
LLama2 13B-hf-chat 78.54 42.0 27.17 61.5
Mistral-7B Instruct-v0.2 CoT 79.33 44.49 27.35 66.5 (-1.04)
78.74 41.71 27.95 65.1
Mistral-8x7B Instruct-v0.1 CoT 90.16 68.61 45.67 77.9 (+0.21)
89.96 70.74 44.31 78.2
GPT-3.5 text-davinci -003 CoT 87.40 50.19 38.19 72.1 (+4.71)
90.35 60.12 40.55 75.7
PoT 95.67 63.84 29.13 78.3 (+1.27)
96.46 64.75 29.53 81.3
ChatGPT gpt-3.5- turbo-0125 CoT 96.46 82.56 58.66 78.6 (+0.57)
97.05 82.03 56.3 83.2
PoT 96.46 77.48 47.64 81.4 (+1.33)
96.06 79.68 47.64 84.9
Table 3: Accuracies (x100) of proposed PEP against zero-shot CoT and PoT.
@@
Model AnsType PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT+SC 87.99 52.31 39.76 71.7 62.94
@@
GPT-3.5 text-davinci -003 (+1.58) (+8.80) (+0.00) (1.90) (+3.07)
ChatGPT gpt-3.5- turbo-0125 CoT+SC 97.83 87.57 67.72 86.3 84.85
98.23 88.48 67.32 88.7 85.68
(+0.40) (+0.91) (-0.40) (+2.40) (+0.83)
Table 4: Solve rates (x100) of proposed PEP against self-consistency CoT. The temperature is 0.7 and each answer are voted from 20 samples.
The main results are presented in Table 3 and Table 4, showing proposed PEP’s performances on various LMs, answer types and decoding strategies. We only test PoT on two GPT models, as its templates might not be designed for these smaller LMs. Considering the token cost, we validate the self-consistency CoT on two GPT models. The comparison and integration with other problem-related methods are shown in table 5, using greedy decoding and turbo-0125, while verifying two different few-shots settings.
PEP performs well in a variety of situations.
Overall, PEP outperforms the standard CoT in most cases, with the exception of Mistral-7B, but demonstrating improvements for Mistral-8x7B. The performance in PoT and self-consistency settings further validates PEP’s effectiveness and adaptability. The enhancement in PoT could be attributed to PEP simplifying parsing difficulties, thereby facilitating code generation (Jiang et al., 2023b). It is also noteworthy that PEP performs remarkably well on davinci, achieving improvements of 9.93% and 8.80% in greedy search and self-consistency, respectively.
PEP can be effectively integrated with other prompting methods.
Unlike the mentioned problem-related methods, PEP aims to enhance the original problem, making it compatible for integration. As shown in Table 5, incorporating PEP improves the performance of these problem-related methods in both k=1 and k=4 few-shot settings, while the combination with the standard CoT also ranks highly in performance. Besides, we observed Self-ask underperforms when k=1. It’s likely because one example might fail to elicit the dynamic QA process and LMs could abruptly terminate after generating follow-up questions.
Dataset
Methods K-shots PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT k=1 97.64 80.59 54.33 81.3 (+0.48)
97.24 79.38 54.72 84.4
L2M k=1 96.85 $\mathbf{82.11}$ 57.09 83.4 (+0.65)
$\mathbf{98.03}$ 80.44 57.09 $\mathbf{84.5}$
PaS* k=1 93.9 77.63 54.33 78.7 (-0.01)
94.09 77.63 55.51 77.3
Self-Ask* k=1 89.57 61.11 40.94 74.2 (+7.63)
95.08 59.29 $\mathbf{58.27}$ 83.7
CoT k=4 $\mathbf{98.03}$ 82.94 56.69 82.3 (+0.99)
$\mathbf{98.03}$ $\mathbf{83.93}$ 59.06 82.9
L2M k=4 97.64 81.65 57.87 82.5 (+1.57)
97.44 81.8 $\mathbf{61.02}$ 85.7
PaS* k=4 96.26 79.76 57.48 77.9 (+1.50)
96.85 81.58 57.48 81.5
Self-Ask* k=4 94.49 78.7 53.54 80.0 (+3.20)
95.47 81.43 55.91 $\mathbf{86.7}$
Table 5: Comparison and integration of PEP with problem-related methods. We utilized exemplars from Zhou et al. (2022a) and adjusted them for other methods* (details in Sec B). The best and second-best results are highlighted in bold and underlined, respectively.
4.3 Distraction Problem
One particular challenge of ill-formed problems is known as distraction problem (Shi et al., 2023): irrelevant sentences can distract LMs to generate errors. These sentences can be completely irrelevant or relevant to the problem but should have no impact on inference. We tested PEP using a subset of GSMIC (Shi et al., 2023). Two metrics are utilized: (1) micro accuracy: averaged accuracy per example, and (2) macro accuracy: averaged accuracy per base problem. Norm is the accuracy normalized by scores on base problems, measuring how a method is affected by the distractors.
PEP effectively mitigates the distraction problems.
As shown in Table 6, PEP surpasses CoT and L2M of both zero-shot and one-shot settings, in micro- and macro- metrics, indicating superior performance in addressing such ill-formed issues. Beyond overall accuracy, PEP also exhibits enhanced robustness as evidenced by norm accuracy. From the macro perspective, the improvements and stability of PEP are also remarkable.
PEP performs well when prompted with prior knowledge.
When the model is consciously prompted to ignore irrelevant content for the given problem, referred to as Irr-Inst. (Shi et al., 2023), we observed significant improvements in CoT, L2M and PEP. Despite this, PEP still outperforms the 0-CoT and 1-L2M by achieving larger improvements for most cases. PEP particularly excels in 2-step problems and norm accuracies. However, its performance on macro accuracies is inferior, potentially due to a conflict between Irr-Inst. and the one-shot exemplar.
Method PEP (ours) Micro Accuracy Macro Accuracy
2 st $\geq$2 st Overall Norm 2 st $\geq$2 st Overall Norm
0-CoT 75.4 78.0 76.7 89.08 31.67 40.0 35.0 48.61
84.2 81.8 83.0 92.32 56.67 30.0 46.0 61.33
(+8.8) (+3.8) (+6.3) (+3.2) (+25.0) (-10.0) (+11.0) (+12.7)
1-CoT 84.2 81.2 82.7 90.98 46.67 32.5 41.0 47.67
87.8 81.8 84.8 95.07 56.67 37.5 49.0 61.25
(+3.6) (+0.6) (+2.1) (+4.1) (+10.0) (+5.0) (+8.0) (+13.6)
1-L2M 84.6 80.8 82.7 90.98 51.67 27.5 42.0 51.22
83.2 84.0 83.6 93.62 50.0 32.5 43.0 58.11
(-1.4) (+3.2) (+0.9) (+2.6) (-1.7) (+5.0) (+1.0) (+6.9)
@@ 4.3 Distraction Problem
0-CoT 80.6 85.8 83.2 97.65 58.33 55.0 57.0 72.15
88.2 87.2 87.7 98.76 61.67 50.0 57.0 76.0
(+7.6) (+1.4) (+4.5) (+1.1) (+3.3) (-5.0) (0.0) (+3.8)
1-CoT 87.4 89.4 88.4 94.55 68.33 62.5 66.0 83.54
90.6 83.6 87.1 96.89 71.67 57.5 66.0 78.57
(+3.2) (-5.8) (-1.3) (+2.3) (+3.3) (-5.0) (0.0) (-5.0)
1-L2M 85.8 85.8 85.8 96.3 61.67 50.0 57.0 74.03
88.0 85.0 86.5 96.86 61.67 45.0 55.0 69.62
(+2.2) (-0.8) (+0.7) (+0.6) (0.0) (-5.0) (-2.0) (-4.4)
Table 6: Micro and macro accuracies (×100) on GSMIC-1k. We follow the metrics and instructions suggested in (Shi et al., 2023).
@@
Figure 3: Breakdown accuracies w.r.t. irrelevant sentence factors (T: Topic, RO: Role Overlap, NR: Num. Range). Lower accuracy suggests the model is more sensitive to that factor.
4.4 Ablation Study
Breakdown analysis of distracting factors.
We evaluated the performance of CoT, L2M and PEP under various distracting factors, as depicted in Fig 3. PEP significantly outperforms the basic baselines on almost all factors, both in micro and macro accuracies, indicating its potential benefits for downstream reasoning by aiding in problem context recognition and parsing. When prompted by the Irr-inst., PEP shows consistent improvements for most cases, except for 1-CoT. Overall, PEP demonstrates better improvements in handling out-of-distribution distraction factors, specifically (1) off-topic sentences, (2) non-overlapping role names, and (3) out-of-range numbers. The impact of these factors is more pronounced on the macro- than the micro- metric.
Method GSM8k SVAMP GSMIC Overall
CoT 82.56 78.6 76.7 79.60
PEP 82.03 83.2 83.0 82.67
Dec 82.79 83.6 79.0 82.50
Elu 81.80 83.1 82.2 82.57
ETD 80.14 82.5 81.1 81.59
Table 7: Ablation of PEP components.
@@ 4.4 Ablation Study
We verify two components in PEP: (1) Dec: decomposing only, and (2) Elu: elucidating only, and (3) EtD: elucidating first then decomposing. As shown in Table 7, both Dec and Elu are required for PEP, with performance varying across datasets. On GSM8k and SVAMP, Dec even outperforms PEP, while Elu is more effective on GSMIC. Besides, EtD consistently performed worse, suggesting that the coordination and operating order of components are also crucial for PEP.
4.5 Error Analysis
Example 1 (PEP misleads implicit meaning) Example 2 (PEP discards important information)
Q: An interior design firm offers installation for \$129.00. It includes hanging 4 mirrors, 2 shelves, 1 chandelier, and 10 pictures. They will install additional items for an extra \$15.00 per item. Angela has 6 mirrors and 2 chandeliers and 20 pictures that she needs hung. How much will this cost her? Q: Mrs. Harrington bought 12 boxes of oranges at the market. She gave her mom and her sister 2 boxes of oranges each. Then she kept 1/4 of the oranges and sold the rest. How many oranges did Mrs. Harrington sell if each box contains 20 oranges?
PEP + CoT: PEP + CoT:
S1: What is the base installation fee for the interior design firm? \$129.00. S2: What is included in the base installation fee? 4 mirrors, 2 shelves, 1 chandelier, and 10 pictures. S3: How much does the interior design firm charge for each additional item installed? Extra \$15.00 per item installed. S4: How many additional items does Angela need installed? Angela needs 6 mirrors, 2 chandeliers, and 20 pictures installed, which is a total of 28 additional items. Solution: The cost of the additional items is 28 items x \$15.00 per item = \$420.00. Therefore, the total cost for Angela’s installation is \$129.00 + \$420.00 = \$549.00. S1: How many boxes of oranges did Mrs. Harrington buy at the market? 12 boxes. S2: How many boxes of oranges did Mrs. Harrington give to her mom and her sister? 2 boxes. S3: How many boxes of oranges did Mrs. Harrington keep for herself? Mrs. Harrington kept 1/4 of the oranges.S4: How many boxes of oranges did Mrs. Harrington sell? The rest. Solution: 1. Calculate the total number of oranges Mrs. Harrington bought: 12x20=240. 2. Calculate the total number of oranges she gave to her mom and her sister: 2x2=4, 4x20=80. 3. Total number of oranges kept for herself: 1/4x240=60. 4. Calculate the total number of oranges sold: 240-80-60=100.
Table 8: Error Analysis via turbo. When pre-processing the problems, PEP may changes the underline meaning and discarding information (red) of the original sentence (cyan).